Gangnam Corp is a South Korea-based music entertainment company, but is interested in launching a new line of clothes and sunglasses under the brand name PSY in hopes of becoming an international phenomenon. They have surveyed the clothing industry and have identified Style Inc. as the closest pure-play firm for the proposed PSY division.

Gangnam Corp has a beta of 1.45 and is 95% equity-financed. Style Inc. has an equity beta of 2.18 and a debt-equity ratio of 2.08. The risk-free rate of return is 1.9 percent and the market risk premium is 9.4 percent.

What cost of equity should Gangnam Corp use for its PSY apparel division? Assume taxes are 34% and that Gangnam Corp will maintain its current capital structure for its new apparel division.
Answer #1
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
2.18 = Unlevered Beta*(1+((1-0.34)*(2.08)))
Unlevered Beta = 0.92

E/A = 0.95

D/A = 1-E/A = 1-0.95 = 0.05

D/E = D/A/(1-D/A) = 0.05/0.95 = 0.05263

Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 0.92*(1+((1-0.34)*(0.05263)))
levered beta = 0.95
As per CAPM
expected return = risk-free rate + beta * (Market risk premium)
Expected return% = 1.9 + 0.95 * (9.4)
Expected return% = 10.83